OpenStudy (anonymous):

A box contains twenty-five cards numbered 1 through 25. One card is to be randomly drawn from the box. What is the probability that the card drawn will show a multiple of 5, a multiple of 6, or a multiple of 7? a. 12/25 b.9/25 c.3/25 d.1/25

OpenStudy (anonymous):

@driftracer305

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@adrynicoleb

OpenStudy (driftracer305):

k this is preety easy

OpenStudy (driftracer305):

when u see OR....... just add up all probabilities

OpenStudy (aum):

In the number range1 to 25: Multiples of 5 are: 5, 10, 15, 20 and 25 Multiples of 6 are: 6, 12, 18, 24 Multiples of 7 are: 7, 14, 21 Number of desired outcomes = 5 + 4 + 3 = 12 numbers. The probability of desired outcome = 12 / 25

OpenStudy (driftracer305):

hey..... even a better way........... @aum ......... lol great work..... i added up all the fractions..

OpenStudy (aum):

thanks.